OpenStudy (anonymous):

This season a softball player scores 6 runs more than three times the number of runs she bats in (also known as RBIs). She scores 108 runs this season. How many RBIs did this player have?

OpenStudy (sleepyjess):

Runs = 3(RBI's) + 6 108 = 3(RBI's) + 6 108-6 = 3(RBI's) + 6 - 6 102 = 3(RBI's) 102/3 = RBI's 34 = RBI's

OpenStudy (sleepyjess):

To check your work 3(34) + 6 = 108 102 + 6 = 108 108 = 108

OpenStudy (anonymous):

Thanks

OpenStudy (sleepyjess):

Your welcome!

OpenStudy (anonymous):

you're the best explainer I've ever had